# AlloyASG: Alloy Predicate Code Representation as a Compact Structurally Balanced Graph

### 29 Feb 2024 v1

@@ 1. Introduction
Alloy (alloy, 16) is a declarative, relational logic-based modeling language for describing structural and dynamic properties of a system’s design. Since Alloy is used to verify software system designs (Zave15, 40, 5, 34, 8), and to perform various forms of analyses over the corresponding implementation, including deep static checking (JacksonVaziri00Bugs, 17, 10), systematic testing (MarinovKhurshid01TestEra, 22), data structure repair (ZaeemKhurshidECOOP2010, 39), automated debugging (GopinathETALTACAS2011, 12) and to synthesize security attacks (websecurity, 2, 23, 29), the consequences could be catastrophic if an Alloy model is faulty yet successfully compiles. Several iterative search methods based on ASTs have already been implemented for Alloy model automation and repair (amra, 32, 14, 42, 7). However, any iterative search methods for mutants or AST elemental changes that fix the Alloy code segments require expansive storage and have intractable search spaces due to the exponential growth of the code segment length, limiting their computational efficiencies.
@@
To create a machine learning algorithm that is aware of the logic of the programming language we are generating code in, we need a logical structure-aware, vectorized, machine-and-human-readable, and computationally efficient code representation that improves upon the AST adjacency matrix representation that grows exponentially in size. Some of the representation learning methods (code2seq, 4, 25) train functions that output vectors from the AST, yet the nature of their dependency on the control flow paths in the AST of those general-purpose languages renders their methodologies implausible for a specification language such as Alloy, which is declarative and lacks control flow. Other methods (astnn, 41, 33, 35) try to focus on the raw AST for a universal representation, yet those algorithms still suffer from rapid growth in the size of the AST. Therefore, this paper aims to create a structure that generates a compact representation of the code segment yet still gives the same complete information as the raw AST. Recent work created an Abstract Semantic Graph (ASG) (DuffyASG, 9), which is close to meeting our data representation needs, but the original designation lacks a numerical representation that provides an adjacency or Laplacian matrix that could be used in popular graph machine learning algorithms (gcn1, 24, 11, 18, 31, 27). Therefore, in this paper, we introduce a matrix representation of the ASG with the following objectives:
@@ 2.3. Alloy Grammar
We consider the Alloy grammar (alloy, 16) before constructing the graph representation. The basic blocks under the overall root of an Alloy model, ModelUnit, consists of a set of declarations, such as signatures as basic types, functions as processes, and predicates as logical arguments. Our dataset Alloy4Fun (alloy4fun, 21) covers only simple predicates without references to external libraries; therefore, creating representations for those code segments included in the dataset is an ideal starting point.
A predicate defines a scope that contains an expression while itself is a solid, invariant root for any fixing pairs. We can easily enumerate all possible nodes under an expression (or formula): the finitely-enumerable types of expressions or formulae, plus variable declarations that only occur under the quantifiers of first-order logic. In this paper, we assign them fixed signatures; yet an adaptive signature system could be used in case of an extension of the CSBASG into either more types of nodes or extra optimizations.
@@ 3. CSBASG Construction
Intuitively, to create an ASG following Definition 3, we first need to do a pre-order traversal on the AST to identify and combine the semantically identical nodes. We start by constructing an isomorphism between $\mathcal{V}$ and $N\cup\Sigma$, then rewrite each AST link into an ASG edge. The isomorphism is trivial since AST nodes with the same word are combined into the same node in the ASG. Nonetheless, constructing the ASG edge set by the AST relations $\xi$ requires a cautious step: merging two AST nodes requires clarifying the orders of the visit to avoid confusion. A naive approach that deviates a bit from the ASG definition is to use two numbers for the weight values $w_{ij}$ instead of one, that is, for any of the existing link $x_{2}\in\xi(x_{1})$ in the AST, defining $v_{1}=\sigma(x_{1}),v_{2}=\sigma(x_{2})$. Then, we create an edge between $v_{1},v_{2}$ with a weight-pair $(\omega,t)\in\mathbb{N}\times\mathbb{N}$, specifying that $x_{2}$ is the $\omega$-th child of $x_{1}$ in the order defined for $\xi(x_{1})$, and it is the $t$-th time to visit $v_{1}$ in the program logic.
@@
Algorithm 1 is a high-level process to convert an AST to a CSBASG. It follows the process of an abstract interpretation, where $\vec{t}$ in the recursive part gives a simplified interpretation state tracking the times of node visitations in the pre-order traversal of the AST. In general, with a static signature assignment $\varphi$, the algorithm terminates in $O(|X|)$ since each of the nodes in the AST was visited exactly once and $|\mathcal{V}|<|X|$ strictly. We say an ASG is generated by an AST if the AST was converted to the same ASG with the same signature assignment and magnitude mapping.
@@
2:Output: a CSBASG $\mathcal{G}(A)$ representing $T$ with its adjacency matrix $A$, and the signatures of the nodes $\{\theta_{1},\theta_{2},...,\theta_{|\mathcal{V}|}\}$, each corresponds to a column of $A$.
@@
5:for $i\leftarrow 2,...,|X|$ do
6:     $node\leftarrow X_{i}$
7:     $\theta_{i}\leftarrow\varphi(\sigma(node))$
8:     if $\theta_{i}\notin\vec{\theta}$ then
9:         $\vec{\theta}\leftarrow\vec{\theta}::\theta_{i}$
10:     end if
11:     $nodemap\leftarrow nodemap::(node,\theta_{i})$
12:end for
13:$A_{\vec{0}}\leftarrow[0]_{|\mathcal{V}|\times|\mathcal{V}|}$
14:$A\leftarrow$ encode-recursive($r,\vec{0}_{|\mathcal{V}|},A_{\vec{0}}$)
15:return $A,\vec{\theta}$
16:function encode-recursive($r_{t},\vec{t},A_{\vec{t}}$)
17:     $\theta_{t}\leftarrow nodemap.\textsc{get}(r_{t})$
18:     $i\leftarrow\vec{\theta}.\textsc{get-index}(\theta_{t})$
19:     $t\leftarrow\vec{t}.\textsc{get}(i)$
20:     $\vec{t}^{\prime}\leftarrow\vec{t}$
21:     $\vec{t}^{\prime}[i]\leftarrow t+1$
22:     $A_{\vec{t}^{\prime}}\leftarrow A_{\vec{t}}$
23:     for $\omega\leftarrow 1,...,|\xi(r_{t})|$ do
24:         $c\leftarrow\xi(r_{t}).\textsc{get}(\omega)$
25:         $\theta_{c}\leftarrow nodemap.\textsc{get}(c)$
26:         $j\leftarrow\vec{\theta}.\textsc{get-index}(\theta_{c})$
27:         $A_{\vec{t}^{\prime}}[j,i]\leftarrow\beta(\mathcal{M}(\omega,t),A_{\vec{t}}[j,i])\cdot\angle(\theta_{i}-\theta_{j})$
28:         $A_{\vec{t}^{\prime}}\leftarrow\textsc{encode-recursive}(c,\vec{t}^{\prime},A_{\vec{t}^{\prime}})$
29:     end for
30:     return $A_{\vec{t}^{\prime}}$
31:end function
@@ 4. Application of CSBASG on Alloy
Figure 1. An example illustrating the compactness and simplicity of CSBASG. The predicate code (upper left) was parsed into its tree form as the AST (upper right), which has 43 nodes, and it could be simplified to the ASG (lower left) with 15 nodes. The table (lower right) is the correspondence of the nodes in the ASG and its description in the code, and so are the signatures defined and multiplied to the polynomial-based encoding values as shown in the graph. Such constructs a CSBASG.
@@
Figure 1 gives an example of the encoding from an Alloy predicate AST to its corresponding CSBASG using a polynomial-based static scheme. The syntactic-semantic pairs in the table give the properties of each node in the AST and combine the AST nodes with identical semantic and syntactic properties. The angular signature assigned to each node is $\Psi D+\psi\delta$, where $\delta=D/128$ for expression or formula nodes and $\delta=D/65536$ for the relational declarations that would appear below a quantifying expression or formula, since by Alloy syntax multiple variables could be declared under the same relational declaration node.
The complex values could be used to compare two ASGs with shared parts and will be discussed in Section 5.3.
@@ 5. Experiment
5.1. Dataset
Our dataset encompasses 6,601 models from Alloy4Fun(alloy4fun, 21), sourced from 8 different categories of problem, filtered from a larger dataset to ensure each is compliable and runnable. Each model has two predicates under a set of definitions of relations; a student writes one named InvX (or PropX, X is an identifying number) and is potentially faulty as a mutant predicate, and another called InvXC is the ground truth that gives a correct formula. Then we run two predicates, overconstrained and underconstrained, which determine if there are cases that satisfy InvX but not InvXC, or vice versa. If a model has no cases for both overconstrained and underconstrained predicates, then InvX is formally correct. This gives four categories of the models: correct ($InvX=InvXC$), overconstrained only ($InvX\subsetneq InvXC$), underconstrained ($InvXC\subsetneq InvX$), and both overconstrained and underconstrained ($InvX\not\subset InvXC\land InvXC\not\subset InvX$), covers all possible cases.
5.2. Metric 1: Compactness of representation
We evaluate the compactness of the representation by the percentage of nodes in the ASG representation that are less than the raw AST. Overall, we achieved an impressive 26.6% reduction of nodes over the dataset, considering there is no loss of information, and the performance for each category of predicates is given in Table 1.
@@ 5.2. Metric 1: Compactness of representation
cv_v1 55 UNDER 31.45 38.05
train 102 CORRECT 24.77 17.4
@@
train 178 UNDER 31.96 24.53
trash_rl 340 CORRECT 16.45 6.43
trash_rl 150 UNDER 13.42 7.1
@@
trash_rl 63 UNDER 31.16 11.5
cv_v2 34 CORRECT 30.74 39.42
@@ 5.3. Metric 2: A case study of comparability
Since the dataset was written by students who have little experience in Alloy modeling and some of them write models that have few, if not zero, common edges; however, among them, there could also be some manual repairs that could be shown as a process of applying a set of “Atomic Mutations”. We call an atomic mutation either adding or removing an edge in the ASG; a relationship change between two nodes could also be represented as removing the old edge and adding the new one.
Figure 2 gives an example of the decoded CSBASG construction for comparing two predicates in the same declarative environment, the faulty and correct solution to the same problem. The common parts, as reflected by the subgraph characterized with black edges, show the correct declaration of the first two variables by the student given the right part of the graph and the trivial correctness of the fields associated with the variables as elements of the pre-declared sets were shown in the bottom-left corner. However, most of the graph shows significant deviation of logic, as expected for the student-written mutant being both overconstrained and underconstrained. Straightforwardly, each of the red or blue edges in the graph is an atomic mutation between the pair of predicates, the red edge is a link to be broken, and the blue edge is a link to be established in a hypothetical automatic fixing application.

### 21 Mar 2024 v3

@@ 1. Introduction
Alloy (alloy, 16) is a declarative, relational logic-based modeling language for describing structural and dynamic properties of a system’s design. Since Alloy is used to verify software system designs (Zave15, 40, 5, 34, 8), and to perform various forms of analyses over the corresponding implementation, including deep static checking (JacksonVaziri00Bugs, 17, 10), systematic testing (MarinovKhurshid01TestEra, 22), data structure repair (ZaeemKhurshidECOOP2010, 39), automated debugging (GopinathETALTACAS2011, 12) and to synthesize security attacks (websecurity, 2, 23, 29), the consequences could be catastrophic if an Alloy model is faulty yet successfully compiles. Several iterative search methods based on ASTs have already been implemented for automated repair and partial generation of Alloy models (amra, 32, 14, 42, 7). However, any iterative search methods for mutants or AST elemental changes that fix the Alloy code segments require expansive storage and have intractable search spaces due to the exponential growth of the code segment length, limiting their computational efficiencies.
@@
To create a machine learning algorithm that is aware of the logic of the programming language we are generating code in, we need a logical structure-aware, vectorized, machine-and-human-readable, and computationally efficient code representation that improves upon the AST adjacency matrix representation that grows exponentially in size. Some of the representation learning methods (code2seq, 4, 25) train functions that output vectors from the AST, yet the nature of their dependency on the control flow paths in the AST of those general-purpose languages renders their methodologies implausible for a specification language such as Alloy, which is declarative and lacks control flow. Other methods (astnn, 41, 33, 35) try to focus on the raw AST for a universal representation, yet those algorithms still suffer from rapid growth in the size of the AST. Therefore, this paper aims to create a structure that generates a compact representation of the code segment, yet still gives the same complete information as the raw AST. Recent work created an Abstract Semantic Graph (ASG) (DuffyASG, 9), which is close to meeting our data representation needs, but the original designation lacks a numerical representation that provides an adjacency or Laplacian matrix that could be used in popular graph machine learning algorithms (gcn1, 24, 11, 18, 31, 27). Therefore, in this paper, we introduce a matrix representation of the ASG with the following objectives:
@@
Therefore, in this paper, we make the following contributions:
CSBASG: We introduce the concept of a CSBASG that meets our three objects (one-to-one correspondence, more compact, and supports direct comparisons). We provide an algorithm to convert an AST into a CSBASG and convert a CSBASG back to an AST and corresponding code segments.
Application to Alloy: We tailor the generation of CSBASG to the Alloy modeling language.
Evaluation: We evaluate the improvement on compactness and density of a CSBASG compared to an AST and evaluate how to compare multiple CSBASGs.
Open Source: We release our source code to produce https://github.com/AlexandervonWu/AlloyASG-Release.
@@ 2. Background
In this section, we describe key components of Abstract Semantic Graphs and the Alloy modeling language.
@@ 2.2. Complex-weighted Graph and its Structural Balance
(a) (b) (c)
1. sig Person  { 2.   Tutors : set Person,3.   Teaches : set Class4. }5. sig Group 6. sig Class  { Groups : Person -> Group }7. sig Teacher in Person  {}8. sig Student in Person  {} /* Assuming a universe of 3 persons, the tutoring  * chain of every person eventually reaches a Teacher. */ 9. pred inv15oracle {10.   all p:Person | some Teacher&(^Tutors).p11. } 12. pred inv15 { 13.   all p : Person { some t : Teacher {14.     t in p.Tutors15.     or t in p.Tutors.Tutors16.     or t in p.Tutors.Tutors.Tutors17.   }}18.}
Figure 1. Alloy Model of a Classroom Management System with an oracle and student submission for inv15
2.3. Alloy
Alloy users write models that describe the properties of the system of interest. Then, the Analyzer helps the user understand their system by displaying the consequences of their properties, helping identify any missing or incorrect properties, and exploring the impact of modifications to those properties. To achieve this, the Analyzer uses off-the-shelf SAT solvers to search for scenarios, which are assignments to the sets and relations of the model such that all executed formulas hold. These scenarios are presented visually and introduce a separation of concern between writing and exploring a specification. If no such scenario can be found, the Analyzer reports that the formulas are unsatisfiable.
To highlight how modeling in Alloy works, Figure 1 depicts the base model of a classroom management system. This model is from the Alloy4Fun dataset, which is a collection of submissions made by novice users learning Alloy (alloy4fun, 21). Signature paragraphs introduce named sets and can define relations, which outline relationships between elements of sets. Line 1 introduces a named set Person and establishes that each Person atom connects to any number of (set) Person atoms through the Tutor relation and each Person atom connects to any number of (set) Class atoms through the Teaches relation. Line 5 introduces the named set Group, which contains no relations. Line 6 introduces the named set Class and states that each class has a set of people assigned to a group using the ternary relational Groups. Lines 7 and 8 introduce the named sets Teacher and Student as subsets (in) of Person.
Predicate paragraphs introduce named first-order, linear temporal logic formulas that can be invoked elsewhere. Figure 1 (b) depicts the oracle for exercise inv15. The predicate inv15oracle uses universal quantification (‘all’), set multiplicity (‘some’), set intersection (‘&’), transitive closure (‘^’), and relational join (‘.’) to express that for every person, there is an intersection between the set Teacher and the set of Tutors reachable by one or more invocation of Tutor’s relation. The student submission in Figure 1 (c) is faulty, and and is overconstrained, which means the faulty formula prevents instances that should have been generated, and underconstrained, which means the faulty formula allows both instances that should have been prevented. The student attempted to use nested quantification to explicitly outline that a teacher should be reachable in 1 to 3 traversals down the Tutor’s relation.
@@ 2.4. Alloy Grammar
Figure 2. An instance of an Alloy predicate in the dataset parsed with pretty string and its raw AST representation.
We consider the Alloy grammar (alloy, 16) before constructing the graph representation. The basic blocks under the overall root of an Alloy model, ModelUnit, consists of a set of declarations, such as signatures as basic types, functions as processes, and predicates as logical arguments. Our dataset Alloy4Fun (alloy4fun, 21) covers only simple predicates without references to external libraries; therefore, creating representations for those code segments included in the dataset is an ideal starting point. Figure 2 gives an example of an Alloy predicate in the dataset that contains 32 AST nodes. From this representation, it is obvious that the raw AST representation has been relatively huge and complicated, even with a short segment of code enveloped in a single predicate, mostly due to the redundancy of the semantically identical nodes presenting in the AST for their multiple appearances in the original code. In the given example, there are three binary operators ‘‘in’’, and 6 dots (JOIN) as operators; these would all be treated as distinct nodes in a raw AST representation.
A predicate defines a scope that contains an expression while itself is a solid, invariant root for any fixing pairs. We can easily enumerate all possible nodes under an expression (or formula): the finitely enumerable types of expressions or formulae, plus variable declarations that only occur under the quantifiers of first-order logic. In this paper, we assign them fixed signatures, yet an adaptive signature system could be used in case of an extension of the CSBASG into either more types of nodes or extra optimizations.
@@ 3. CSBASG Construction
Intuitively, to create an ASG following Definition 3, we first need to do a pre-order traversal on the AST to identify and combine the semantically identical nodes. We start by constructing an isomorphism between $\mathcal{V}$ and $N\cup\Sigma$, then rewrite each AST link into an ASG edge. The isomorphism is trivial since AST nodes with the same word are combined into the same node in the ASG. Nonetheless, constructing the ASG edge set by the AST relations $\xi$ requires a cautious step: merging two AST nodes requires clarifying the orders of the visit to avoid confusion. or instance, in the example given in Figure 1, if we consider the binary operator ‘‘IN’’ as a single node without specifying the order of execution, we would not specify which could be put at the right of the operator, ‘‘p.Tutors’’, ‘‘p.Tutors.Tutors’’ or ‘‘p.Tutors.Tutors.Tutors’’. A naive approach that deviates a bit from the ASG definition is to use two numbers for the weight values $w_{ij}$ instead of one, that is, for any of the existing link $x_{2}\in\xi(x_{1})$ in the AST, defining $v_{1}=\sigma(x_{1}),v_{2}=\sigma(x_{2})$. Then, we create an edge between $v_{1},v_{2}$ with a weight-pair $(\omega,t)\in\mathbb{N}\times\mathbb{N}$, specifying that $x_{2}$ is the $\omega$-th child of $x_{1}$ in the order defined for $\xi(x_{1})$, and it is the $t$-th time to visit $v_{1}$ in the program logic.
@@
Algorithm 1 is a high-level process to convert an AST to a CSBASG. It follows the process of an abstract interpretation, where $\vec{t}$ in the recursive part gives a simplified interpretation state tracking the times of node visitations in the pre-order traversal of the AST. We initialize the nodemap and the list of angular signatures for each node, with the only element being the dummy overall root, and assign each node an angular signature with regard to the predefined assignment, outlining every unique signature; the number of unique angular signatures (semantically distinct nodes) is the number of rows and columns of the resulting adjacency matrix. After that, we initialize a state vector $\vec{t}$ to count the times of the order of the visit to each node. A recursive function will then iterate over the AST beginning from the overall root, update the adjacency matrix for each downlink from the root node according to the composition function $\beta$ and the encoding method $\mathcal{M}$, while updating the counter of the given node with the corresponding entry of the state vector signaling the pre-order time of identical appearance in the AST before recursively visit the subtree with the new root being a child of the local root.
In general, with a static signature assignment $\varphi$, the algorithm terminates in $O(|X|)$ since each of the nodes in the AST was visited exactly once and $|\mathcal{V}|<|X|$ strictly. We say an ASG is generated by an AST if the AST was converted to the same ASG with the same signature assignment and magnitude mapping.
@@
2:Output: a CSBASG $\mathcal{G}(A)$ representing $T$ with its adjacency matrix $A$, and the signatures of the nodes $\{\theta_{0},\theta_{1},\theta_{2},...,\theta_{|\mathcal{V}|}\}$, each corresponds to a column of $A$.
@@
5:$\mathcal{V}\leftarrow\{r\}$
6:for $i\leftarrow 2,...,|X|$ do
7:    $node\leftarrow X_{i}$
8:    $\theta_{i}\leftarrow\varphi(\sigma(node))$
9:    if $\theta_{i}\notin\vec{\theta}$ then
10:         $\vec{\theta}\leftarrow\vec{\theta}::\theta_{i}$
11:         $\mathcal{V}\leftarrow\mathcal{V}::node$
12:    end if
13:    $nodemap\leftarrow nodemap::(node,\theta_{i})$
14:end for
15:$A_{\vec{0}}\leftarrow[0]_{|\mathcal{V}|\times|\mathcal{V}|}$
16:$A\leftarrow$ encode-recursive($r,\vec{0}_{|\mathcal{V}|},A_{\vec{0}}$)
17:return $A,\vec{\theta}$
18:function encode-recursive($r_{t},\vec{t},A_{\vec{t}}$)
19:    $\theta_{t}\leftarrow nodemap.\textsc{get}(r_{t})$
20:    $i\leftarrow\vec{\theta}.\textsc{get-index}(\theta_{t})$
21:    $t\leftarrow\vec{t}.\textsc{get}(i)$
22:    $\vec{t}^{\prime}\leftarrow\vec{t}$
23:    $\vec{t}^{\prime}[i]\leftarrow t+1$
24:    $A_{\vec{t}^{\prime}}\leftarrow A_{\vec{t}}$
25:    for $\omega\leftarrow 1,...,|\xi(r_{t})|$ do
26:         $c\leftarrow\xi(r_{t}).\textsc{get}(\omega)$
27:         $\theta_{c}\leftarrow nodemap.\textsc{get}(c)$
28:         $j\leftarrow\vec{\theta}.\textsc{get-index}(\theta_{c})$
29:         $A_{\vec{t}^{\prime}}[j,i]\leftarrow\beta(\mathcal{M}(\omega,t),A_{\vec{t}}[j,i])\cdot\angle(\theta_{i}-\theta_{j})$
30:         $A_{\vec{t}^{\prime}}\leftarrow\textsc{encode-recursive}(c,\vec{t}^{\prime},A_{\vec{t}^{\prime}})$
31:    end for
32:    return $A_{\vec{t}^{\prime}}$
33:end function
@@ 4. Application of CSBASG on Alloy
Figure 3. An example illustrating the compactness and simplicity of CSBASG; the original code (upper) could be expressed as an equivalent multigraph (middle) with the AST in Figure 1, and turns into the adjacency relationship (lower) with the polynomial-based encoding. With the same predicate in Figure 1, the number of nodes reduced significantly from 31 to 12. All edges are now assigned an integer weight value, which forms an adjacency matrix.
@@
Figure 3 gives an example of the encoding from an Alloy predicate AST to its corresponding CSBASG using a polynomial-based static scheme. The syntactic-semantic pairs in the table give the properties of each node in the AST and combine the AST nodes with identical semantic and syntactic properties. The angular signature assigned to each node is $\Psi D+\psi\delta$, where $\delta=D/128$ for expression or formula nodes and $\delta=D/65536$ for the relational declarations that would appear below a quantifying expression or formula, since by Alloy syntax multiple variables could be declared under the same relational declaration node. In our example, there are six JOIN binary operators (.), which are represented as six distinct nodes in the original AST; in the ASG approach, the nodes are combined as a single node with the label BinaryExpr.JOIN.
The complex values could be used to compare two ASGs with shared parts, which will be discussed in Section 5.3.
@@ 5. Experiment
cv_v1 54 UNDER 32.03 37.94
train 75 CORRECT 29.44 19.43
@@
train 174 UNDER 32.70 24.43
trash_rl 252 CORRECT 20.25 6.73
trash_rl 150 UNDER 13.42 7.10
@@
trash_rl 62 UNDER 31.67 11.68
cv_v2 31 CORRECT 30.95 40.47
@@
5.1. Dataset
Our dataset encompasses 6,307 models from Alloy4Fun(alloy4fun, 21), sourced from 8 different categories of problem, filtered from a larger dataset to ensure each is compliable, runnable, nonempty, and the student-written model is not completely identical with the ground truth. Each model has two predicates under a set of definitions of relations; a student writes one named InvX (or PropX, X is an identifying number) and is potentially faulty as a mutant predicate, and another called InvXC is the ground truth that gives a correct formula. Then we run two predicates, overconstrained and underconstrained, which determine if there are cases that satisfy InvX but not InvXC, or vice versa. If a model has no cases for both overconstrained and underconstrained predicates, then InvX is formally correct. This gives four categories of the models: correct ($InvX=InvXC$), overconstrained only ($InvX\subsetneq InvXC$), underconstrained ($InvXC\subsetneq InvX$), and both overconstrained and underconstrained ($InvX\not\subset InvXC\land InvXC\not\subset InvX$), covers all possible cases.
5.2. Metric 1: Compactness of representation
We evaluate the compactness of the representation by the percentage of nodes in the ASG representation that are less than the raw AST. Overall, we achieved an impressive 27.25% reduction of nodes over the dataset, considering there is no loss of information, and the performance for each category of predicates is given in Table 1.
@@ 5.3. Metric 2: A case study of comparability
Since the dataset was written by students who have little experience in Alloy modeling and some of them write models that have few, if not zero, common edges; however, among them, there could also be some manual repairs that could be shown as a process of applying a set of ‘‘Atomic Mutations’’. We call an atomic mutation either adding or removing an edge in the ASG; a relationship change between two nodes could also be represented as removing the old edge and adding the new one.
Figure 4 gives an example of the decoded CSBASG construction for comparing two predicates in the same declarative environment, the faulty and correct solution to the same problem. The common parts, as reflected by the subgraph characterized with black edges, show the correct declaration of the first two variables by the student given the right part of the graph and the trivial correctness of the fields associated with the variables as elements of the pre-declared sets were shown in the bottom-left corner. However, most of the graph shows significant deviation of logic, as expected for the student-written mutant being both overconstrained and underconstrained. Straightforwardly, each of the red or blue edges in the graph is an atomic mutation between the pair of predicates; the red edge is a link to be broken, and the blue edge is a link to be established in a hypothetical automatic fixing application. In the example above, we could see that out of 41 edges shown in Figure 4, there are 10 common edges presenting in both the faulty model and the ground truth, 20 edges are added and 11 edges are removed from the faulty model, which are performs of atomic mutations.
[…]
Dataset #Models Mut.Type MutantCorrect%
lts 61 CORRECT 51.61 49.31
lts 113 UNDER 69.66 31.03
lts 138 BOTH 71.46 36.55
lts 66 UNDER 60.87 42.91
classroom_fol 495 CORRECT 82.07 71.77
classroom_fol 223 UNDER 77.54 63.51
classroom_fol 1115 BOTH 79.21 64.01
classroom_fol 166 UNDER 69.61 61.49
cv_v1 52 CORRECT 62.88 38.86
cv_v1 118 UNDER 58.99 34.84
cv_v1 68 BOTH 66.96 36.03
cv_v1 54 UNDER 69.96 37.52
train 75 CORRECT 66.66 63.86
train 170 UNDER 70.06 49.00
train 277 BOTH 75.16 50.97
train 174 UNDER 73.29 60.55
trash_rl 252 CORRECT 90.64 73.49
trash_rl 150 UNDER 84.08 61.93
trash_rl 267 BOTH 89.39 62.35
trash_rl 62 UNDER 90.21 75.24
cv_v2 31 CORRECT 63.93 38.38
cv_v2 40 UNDER 65.78 33.73
cv_v2 21 BOTH 67.67 35.04
cv_v2 12 UNDER 70.70 36.82
classroom_rl 495 CORRECT 82.07 71.77
classroom_rl 223 UNDER 77.54 63.51
classroom_rl 1115 BOTH 79.21 64.01
classroom_rl 166 UNDER 69.61 61.49
production 25 CORRECT 68.12 49.40
production 25 UNDER 71.80 45.60
production 22 BOTH 70.67 42.82
production 36 UNDER 43.86 45.62
Table 2. Comparability quantification results in the per-category experiment. The percentages given are the ratio of the edges that also presenting in its mutant or corrected counterpart with the same relationship (position and execution order), that is the percentage of the shared edges in each predicate.